### v2

@@ Abstract
A self-consistent model is developed to explain the anomalously large enrichment of nickel isotopes observed in ablation plumes from ultrafast laser irradiation of solid surfaces. The model is based on the spontaneous creation of a magnetic centrifuge in the ablation plume and the associated cyclotron rotation of plasma ions with rotation rates on the order of $10^{9}$ radians per second. Mass separation occurs around the radial coordinate of cylindrical symmetry with longitudinal axis normal to the ablating surface. A Gaussian shaped radial magnetic field distribution is extracted for Ni isotopes with peak field of 53 Megagauss and average field of 20 Megagauss. In addition to cyclotron rotation of ions, a rigid rotor model is also presented that is associated with the hydrodynamic rotation of the entire plasma and is shown to be of little consequence for the isotope enrichment. Cyclotron rotations dominate the process. A profound resonance of enrichment is observed for specific cyclotron charge states and is shown to be associated with Ion Bernstein Wave (IBW) coupling that provides electrostatic acceleration to these resonant orbits.
@@ 1 Introduction
In a paper published in 1999 we reported unusually strong isotope enrichment in the normal ejection direction from ultrafast femtosecond laser ablation pulses using laser power densities in the $10^{13}-10^{15}$ W/cm2 Pronko et al. (1999). It was postulated, at that time, based on the angular distribution of those enrichment and concentration distributions of deposited material, that an intense Megagauss magnetic field was responsible for the observed effects. Since then, a number of papers have appeared in the literature reporting on experimental observations of such fields Kahaly et al. (2009); Pisarczyk et al. (2015); Sandhu et al. (2002, 2001) . In view of these developments and the ongoing interest in such magnetic fields, we are revisiting our earlier work and providing important additional data that was obtained during that period, and that has hitherto not been published but which bear directly on the subjectVan Rompay (2003). At the time of that earlier 1999 paper, we were criticized for proposing a magnetic field effect for our observed results, with alternative explanations being proposed using hydrodynamic separation as being responsible for the observed phenomena Gupta and Naik (2001). This criticism was leveled at our work, at that time, out of an incomplete understanding of the characteristics of ultrafast ablation plumes and the various time-dependent processes that occur during its temporal development. In addition, thermodynamic arguments that are relevant to nanosecond laser ablation were being referenced and used in those criticisms. The difference between femtosecond and nanosecond laser ablation pulses is now known to be associated with the fact that femtosecond ablation plumes are characterized by, among other things, a Coulomb explosion emanating out of a solid density plasma VanRompay et al. (1998) and not by the thermodynamic plasma formation associated with nanosecond ablation plasmas.
In the present paper, we will present data showing that the elemental and isotopic spatial separations occurring in ultrafast laser ablation plumes can be consistently described by a magnetic centrifuge process occurring in the self-generated magnetic fields of these ultrafast ablation plasmas. Further we shall show that the magnetic centrifuge can be described by specific Larmor orbits of ions within the plasma with possible assistance from an ExB electric field driven bulk plasma rotation. Details concerning resonant separation enhancement of specific charge states by internal wave excitation will also be presented and discussed.
@@ 1.1 Ablation Dynamics
The generation of a plasma from a solid surface by a laser pulse involves the absorption of photon energy by the electrons during the time of the laser pulse and the transfer of that energy to the atoms and ions in the material structure. This transfer of energy to the atoms ultimately results in an ablation plume emanating from the surface. The process of that energy transfer is specific to the laser pulse characteristics and the internal dynamics of the plasma formation. It is important to recognize here that the ablation plume, which is emitted normal to the target surface, will have both spacial and temporal characteristics. We examine the spacial distribution of ions in this report but need to recognize that the plume develops not only with ions, but also with neutrals and nano-particles. This occurs along a specific time line where the pulse absorption is characterized initially by a solid density plasma with fast and slow ions followed by neutrals, and subsequently followed by a burst of nano-particles Albert et al. (2003). These events are observed to have individual time horizons and can, in principle, be separated and studied using mechanical or electrostatic separation mechanisms. Harvesting isotopic ions lends itself to some form of electrostatic separation which will be discussed at a later time.
@@
In regard to the initial plasma state, it is considered axiomatic that a plasma will self-organize after being formed as it relaxes towards a quasi- equilibrium transitional state before eventually dispersing Boyd and Sanderson (2003); Bittencourt (2013). The specific ways this happens will depend on factors such as the electro-magnetic field details that are present and the characteristics of the plasma itself. In the case of high intensity ultrafast laser pulses, as used in this work, the pulse time scales are femtoseconds and intensities on the order of $10^{14}$ to $10^{15}$ watts/cm2. Due to the short time scales and the rapid absorption of energy by the lighter electrons, a coulomb explosion occurs at the surface whereby hot electrons which are ejected outward create a charge imbalance leaving a positive surface layer of ions that are accelerated outward giving them energies in the keV range (see Appendix 8.1) VanRompay et al. (1998). This explosion and transfer of energy happens over a time scale on the order of a few picoseconds or less. During this time interval, electrons also circulate above the target surface and along the low density plasma sheath, acting together with internal thermoelectric currents to form toroidal and longitudinal magnetic fields within the plasma.
@@
The expansion of the initial solid density plasma can be considered to evolve as a one-dimensional (1-D) process, moving as an expanding plug along the normal to the surface of the ablating material Puell et al. (1970). The conversion to a three-dimensional (3-D) expansion occurs after a distance that is on the order of or somewhat greater than the laser beam spot size Stevefelt and Collins (1991). Within the one-dimensional region of the expanding plasma, both toroidal and longitudinal magnetic fields of very high intensity have been observed and measured Kahaly et al. (2009); Pisarczyk et al. (2015); Sandhu et al. (2002, 2001).
@@
The toroidal, or azimuthal, field is considered to form first from rapid outflow of hot electrons under the influence of such mechanisms as the Biermann battery effect. Subsequently a longitudinal field along the one-dimensional z axis forms due to a dynamo effect from a radially expanding, slightly rotating plume that will naturally convert some of a strong toroidal field $B_{\theta}$ into an axial (longitudinal) field $B_{z}$. This occurs primarily through the MHD induction term $\nabla\times(\mathbf{v}\times\mathbf{B})$ together with turbulence such as rotor rotation, tangential surface loops, and non-symmetric three dimensional charge expansion to establish, within picoseconds, a well defined longitudinal field Briand et al. (1985). In addition, the toroidal field will continue to act producing a magnetic pinch on the ionic component of the plasma Yutong et al. (2000); Jackson (2021). This pinching effect will work to hold the plasma together within the one-dimensional zone. The longitudinal magnetic field induces a rotational motion to the expanding plasma and to the energetic particles within, occurring through transverse velocity components and drift rotation, similar to the way it occurs in a conventional plasma centrifuge device Krishnan et al. (1981). The combined effect of these two phenomena, pinching and rotation, can be expected to induce a radial separation of ionic masses within the 1-D cylindrical region of the expansion zone. To the extent that the distribution of ions in the cylindrical zone reaches some quasi equilibrium state, it is expected that these rotational processes will achieve a centrifuge type distribution Krishnan et al. (1981). The plasma behavior in magnetic centrifuge devices has been studied in considerable detail. For the case of an electro-magnetically-driven plasma centrifuge, the rotation rate can be one of two distinct angular frequencies, one at the cyclotron frequency, and one at a specific lower rate, proportional to the $\vec{E}\times\vec{B}$ drift rotation. Derivation and discussion of these two rotational mechanisms are provided in references Kim et al. (1987). In conventional rigid-rotor plasma devices, only the lower of the two characteristic rotation rates is usually observed. This most probably is a consequence of the relatively low magnetic field strengths($\approx 5$kG) used in those devices. In contrast, we show below that our data demonstrates ion rotations at cyclotron frequencies, in these ultrafast ablation plasmas, where much higher field strengths are present (10 -50 MG).
@@ 1.2 Laser Parameters and Detector-Target Geometry
Figure 1: Ablation Model for Enrichment Data as a Function of Observation Angle $\theta$
@@ 2.1 Data
Two different ablation experiments are presented below. An experiment to test the ablative separation of chemical elements has been performed on a Ni/Cu alloy of 45%Ni and 55% Cu. Figure 2 shows the distribution of the Ni/Cu element separation ratios as a function of charge state and angle for the specific angles of 0, 15. 30, and 60 degrees. Charge states +1 through +10 are shown as gray bars, with the highest enrichment labeled. The total enrichment ratio for all ions is plotted as circles. The enrichment ratio is normalized to the element composition ratio. The gray bars represent the contribution of the various charge states from +1 to +10. These data are presented as clusters at the specific angles of 0, 15 ,30, and 60. It is noted that each group has a dominantly enriched charge state along with sub-dominant peaks. These will be discussed further below. Figure 3 shows a similar plot of the $Ni^{58}/Ni^{60}$ isotope ratios, obtained from a pure nickel target, also showing especially strong dominance of particular charge state enrichments. Figure 4 presents ion density distributions as a function of charge state and energy. These ion population data are used to extract a mean kinetic energy per ion to replace the Maxwellian temperature in the ablative plasma. Since the mean kinetic energy of an ideal gas is $(3/2)k_{B}T$ it is appropriate in the present case to replace $2k_{b}T$ with the mean kinetic energy of $1.05$ kV from figure 4 to be used for further computational purposes. Further discussion of Figure 4 is presented in Appendix section 8.1.
@@ 3.1 Analysis of Data
which yields the specific angular rotation rates for each observed angle from Fig. 5.
@@ 3.2 Plasma Rotational Dynamics
In ultrafast laser ablation, the intense laser pulse, which itself has a Gaussian spatial profile Diels and Rudolph (2006), initially creates a highly ionized solid density plasma-plug at the point of impact with the target surface. Modeling this ablation zone as a cylinder with radius on the order-of or somewhat-larger-than the beam spot we can expect to have, within that zone: (a) strong electric fields due to charge separation and coulomb explosion, (b) strong, self-generated magnetic fields due to thermal and charge density gradients and hot electron transport, and (c) rapid plasma expansion transitioning from a dense cylindrical plasma to an expanding conical structure.
Within the initial cylinder there may exist two distinct types of rotational motion. One is the so called ExB drift rotation where the plasma undergoes bulk rotation as a whole. This results from the interaction between radial electric fields and axial magnetic fields. The entire plasma column undergoes rigid-body like rotation with angular rotation rate as given by Chen (2016)
@@
$\omega_{c}=\frac{qB}{m}$ (7)
which depends on mass m, charge q, and relevant magnetic field B leading to species separation, whereas the rigid rotor rotation relies on centrifugal acceleration to accomplish separation. Both modes of separation require a longitudinal magnetic field directed along the z axis of the rotating cylinder.
Equation 7 is found to be critical for the present work. It provides a way of extracting magnetic fields based on cyclotron rotation rates extracted from observed mass separation ratios. Shown below in Figure 6 are the results of using Eq. 7 to calculate the magnetic fields experienced by the alloy elements and the isotopes using their extracted rotation rates as shown in Table 1 above. The numerical average taken from all the data points in Fig.6 provides a field value of 19.3 Megagauss [MG] within the ablation cylinder. These data in Fig. 6 are typical field strengths directly observed experimentally and reported in the literature. Kahaly et al. (2009); Sandhu et al. (2001, 2002, 2004); Forestier-Colleoni et al. (2019); Zhang et al. (2000); Pisarczyk et al. (2015, 2017).
Figure 6: Gaussian and quadratic fit to cyclotron rotation rates and magnetic field as function of plasma radius and observation angle.
@@
Given this information about rotation rates, we can estimate the number of rotations an ion experiences in traveling the 125 $\mu$ length of our idealized centrifuge cylinder. We show, in Table 2 below, the distance traveled per rotation and the number of rotations achieved by an 0.2 keV Ni isotope ion traveling the length of the cylinder. It is seen in Figure 4 that a charge state one ion has this representative energy and represents the bulk of the ionic component in the plasma. The results in Table 2 are computed using separation ratios from Figure 3 and rotation rates from Eq.3. In Fig.3 the total enrichment ratios (for all charge states at a given angle) is shown as circles connected by a solid line. The distance traveled per rotation scales as $E^{1/2}$ for the higher energy ions with a commensurate decrease in number of predicted rotations. At 60 degrees (which has the slowest rotation rate), one full rotation would be accomplished by a 3.2 keV ion traveling the 125 microns. This represents a charge state between 4 and 5 according to Figure 9.

### v4

@@ Abstract
A self-consistent model is developed to explain the anomalously large enrichment of nickel isotopes observed in ablation plumes from ultrafast laser irradiation of solid surfaces. The model is based on the spontaneous creation of a magnetic centrifuge in the ablation plume and the associated cyclotron rotation of plasma ions with effective rotation rates on the order of $10^{9}$ radians per second. Mass separation occurs around the radial coordinate of cylindrical symmetry with longitudinal axis normal to the ablating surface. A Gaussian shaped radial magnetic field $B_{eff}$ is extracted for Ni isotopes which is shown to be a combination of an axial $B_{z}$ component and a second contribution $B_{ibw}$ that represents the equivalent of an “effective magnetic field” contributing to the isotopic separation due to broad spectrum Ion Bernstein Waves providing electrostatic acceleration to the cyclotron orbits. These IBWs are also responsible for a profound resonance of enrichment observed for certain specific charge states. In addition to cyclotron rotation of ions, a rigid rotor model is also presented that is associated with the hydrodynamic rotation of the entire plasma and is shown to be of little consequence for the isotope enrichment. Cyclotron rotation and IBWs dominate the process.
@@ 1 Introduction
In a paper published in 1999 we reported unusually strong isotope enrichment in the normal ejection direction from ultrafast femtosecond laser ablation pulses using laser power densities in the $10^{13}-10^{15}$ W/cm2 pronko1999isotope. It was postulated, at that time, based on the angular distribution of those enrichment and concentration distributions of deposited material, that an intense Megagauss magnetic field was responsible for the observed effects. Since then, a number of papers have appeared in the literature reporting on experimental observations of such fields kahaly2009polarimetric; pisarczyk2015space; pisarczyk2017kinetic; sandhu2002laser; sandhu2001femtosecond . In view of these developments and the ongoing interest in such magnetic fields, we are revisiting our earlier work and providing important additional data that was obtained during that period, and that has hitherto not been published but which bear directly on the subjectvan2003mass. At the time of that earlier 1999 paper, we were criticized for proposing a magnetic field effect for our observed results, with alternative explanations proposing hydrodynamic separation as being responsible for the observed phenomena gupta2001comment. This criticism was leveled at our work, at that time, out of an incomplete understanding of the characteristics of ultrafast ablation plumes and the various time-dependent processes that occur during its temporal development. In addition, thermodynamic arguments that are relevant to nanosecond laser ablation were being referenced and used in those criticisms. The difference between femtosecond and nanosecond laser ablation pulses is now known to be associated with the fact that femtosecond ablation plumes are characterized by, among other things, a Coulomb explosion emanating out of a solid density plasma vanrompay1998pulse and not by the thermodynamic plasma formation associated with nanosecond ablation plasmas. In further support of a centrifuge process it is noted that hydrodynamic separation would show a $1/\sqrt{m}$ mass dependence whereas we observe a mass difference $\Delta m$ dependence as predicted by a centrifuge process vanrompay2000isotope. In the present paper, we will present data showing that the elemental and isotopic spatial separations can be consistently described by a magnetic centrifuge process occurring in the self-generated magnetic fields of these ultrafast ablation plasmas. Further we shall show that the magnetic centrifuge can be described by specific Larmor orbits of ions within the plasma with assistance from resonant and broad spectrum ion Bernstein waves (IBW). Details concerning resonant separation enhancement of specific charge states by these internal wave excitations is presented and discussed. Rigid rotor plasma rotation is demonstrated to be of no significance in the isotope enrichment process.
@@ 1.1 Ablation Dynamics
The generation of a plasma from a solid surface by a laser pulse involves the absorption of photon energy by the electrons during the time of the laser pulse and the transfer of that energy to the atoms and ions in the material structure. This transfer of energy to the atoms ultimately results in an ablation plume emanating from the surface. The process of that energy transfer is specific to the laser pulse characteristics and the internal dynamics of the plasma formation. It is important to recognize here that the ablation plume, which is emitted normal to the target surface, will have both spacial and temporal characteristics. We examine the spacial distribution of ions in this report but need to recognize that the plume develops not only with ions, but also with neutrals and nano-particles. This occurs along a specific time line where the pulse absorption is characterized initially by a solid density plasma with fast and slow ions followed by neutrals, and subsequently followed by a burst of nano-particles albert2003time. These events are observed to have individual time horizons and can, in principle, be separated and studied using mechanical or electrostatic separation mechanisms. Harvesting isotopic ions lends itself to some form of electrostatic separation which will be discussed at a later time.
@@
In regard to the initial plasma state, it is considered axiomatic that a plasma will self-organize after being formed as it relaxes towards a quasi- equilibrium transitional state before eventually dispersing boyd2003physics; bittencourt2013fundamentals. The specific ways this happens will depend on factors such as the electro-magnetic field details that are present and the characteristics of the plasma itself. In the case of high intensity ultrafast laser pulses, as used in this work, the pulse time scales are femtoseconds and intensities on the order of $10^{14}$ to $10^{15}$ watts/cm2. Due to the short time scales and the rapid absorption of energy by the lighter electrons, a coulomb explosion occurs at the surface whereby hot electrons which are ejected outward create a charge imbalance leaving a positive surface layer of ions that are accelerated outward giving them energies in the keV range (see Appendix 8.1) vanrompay1998pulse. This explosion and transfer of energy happens over a time scale on the order of a few picoseconds or less. During this time interval, electrons also circulate above the target surface and along the low density plasma sheath, acting together with internal thermoelectric currents to form toroidal and longitudinal magnetic fields within the plasma.
@@
The expansion of the initial solid density plasma can be considered to evolve as a one-dimensional (1-D) process, moving as an expanding plug along the normal to the surface of the ablating material puell1970heating. The conversion to a three-dimensional (3-D) expansion occurs after a distance that is on the order of or somewhat greater than the laser beam spot size stevefelt1991modelling. Within the one-dimensional region of the expanding plasma, both toroidal and longitudinal magnetic fields of very high intensity have been observed and measured kahaly2009polarimetric; pisarczyk2015space; sandhu2002laser; sandhu2001femtosecond.
@@
The toroidal, or azimuthal, field is considered to form first from rapid outflow of hot electrons under the influence of such mechanisms as the Biermann battery effect. Subsequently a longitudinal field along the one-dimensional z axis forms due to a dynamo effect from a radially expanding, slightly rotating plume that will naturally convert some of a strong toroidal field $B_{\theta}$ into an axial (longitudinal) field $B_{z}$. This occurs primarily through the MHD induction term $\nabla\times(\mathbf{v}\times\mathbf{B})$ together with turbulence such as rotor rotation, tangential surface loops, and non-symmetric three dimensional charge expansion to establish, within picoseconds, a well defined longitudinal field briand1985axial. In addition, the toroidal field will continue to act producing a magnetic pinch on the ionic component of the plasma yutong2000observation; jackson2021classical. This pinching effect will work to hold the plasma together within the one-dimensional zone. The longitudinal magnetic field induces a rotational motion to the expanding plasma and to the energetic particles within, occurring through transverse velocity components and drift rotation, similar to the way it occurs in a conventional plasma centrifuge device krishnan1981plasma. The combined effect of these two phenomena, pinching and rotation, can be expected to induce a radial separation of ionic masses within the 1-D cylindrical region of the expansion zone. To the extent that the distribution of ions in the cylindrical zone reaches some quasi equilibrium state, it is expected that these rotational processes will achieve a centrifuge type distribution krishnan1981plasma. The plasma behavior in magnetic centrifuge devices has been studied in considerable detail. For the case of an electro-magnetically-driven plasma centrifuge, the rotation rate can be one of two distinct angular frequencies, one at the cyclotron frequency, and one at a specific lower rate, proportional to the $\vec{E}\times\vec{B}$ drift rotation. Derivation and discussion of these two rotational mechanisms are provided in references kim1987equilibria; gueroult2018exb. In conventional rigid-rotor plasma devices, only the lower of the two characteristic rotation rates is usually observed. This most probably is a consequence of the relatively low magnetic field strengths($\approx 5$kG) used in those devices. In contrast, we show below that our data demonstrates ion rotations at cyclotron frequencies, in these ultrafast ablation plasmas, where much higher field strengths are present.
@@ 2 Equations and Data
Figure 1: Ablation Model for Enrichment Data as a Function of Observation Angle $\theta$. In the present work $r_{m}$ is taken as 125 microns which is 1.5 times the 50 $\mu\text{m}$ radius of the $(1/e^{2})$ laser intensity profile
@@ 2.1 Data
Two different ablation experiments are presented below. An experiment to test the ablative separation of chemical elements has been performed on a Ni/Cu alloy of 45%Ni and 55% Cu. Figure 2 shows the distribution of the Ni/Cu element separation ratios as a function of charge state and angle for the specific angles of 0, 15. 30, and 60 degrees. Charge states +1 through +10 are shown as gray bars, with the highest enrichment labeled. The total enrichment ratio for all ions is plotted as circles. The enrichment ratio is normalized to the element composition ratio. The gray bars represent the contribution of the various charge states from +1 to +10. These data are presented as clusters at the specific angles of 0, 15 ,30, and 60. It is noted that each group has a dominantly enriched charge state along with sub-dominant peaks. These will be discussed further below. Figure 3 shows a similar plot of the $Ni^{58}/Ni^{60}$ isotope ratios, obtained from a pure nickel target, also showing especially strong dominance of particular charge state enrichments. Figure 4 presents ion density distributions as a function of charge state and energy. These ion population data are used to extract a mean kinetic energy per ion to replace the Maxwellian temperature in the ablative plasma.
@@
Since the mean kinetic energy of an ideal gas is $(3/2)k_{B}T$, it is appropriate in the present case to replace $2k_{b}T$ with the mean kinetic energy of $1.05$ kV from figure 4. Further discussion of Figure 4 is presented in Appendix section 8.1.
@@ 3.1 Analysis of Data
which yields the specific angular rotation rates for each observed angle from Fig. 5.  [^1] It should be noted that $\omega(\theta)$ in Eq. 3 is model dependent since a choice of $r_{\text{max}}$ is required. We have estimated it at $125\,\mu\text{m}$ in all of our subsequent calculations. This choice directly affects the rotation rates extracted from the data. Choosing an $r_{\text{max}}$ of $150\,\mu\text{m}$ would reduce the rotation rates by approximately $17\%$.
@@ 3.2 Plasma Rotational Dynamics
In ultrafast laser ablation, the intense laser pulse, which itself has a Gaussian spatial profile diels2006ultrashort, initially creates a highly ionized solid density plasma-plug at the point of impact with the target surface. Modeling this ablation zone as a cylinder with radius on the order-of or somewhat-larger-than the beam spot we can expect to have, within that zone: (a) strong electric fields due to charge separation and coulomb explosion, (b) strong, self-generated magnetic fields due to thermal and charge density gradients and hot electron transport, and (c) rapid plasma expansion transitioning from a dense cylindrical plasma to an expanding conical structure.
Within the initial cylinder there may exist two distinct types of rotational motion. One is the so called ExB drift rotation where the plasma undergoes bulk rotation as a whole. This results from the interaction between radial electric fields and axial magnetic fields. The entire plasma column undergoes rigid-body like rotation with angular rotation rate as given by chen2016introduction
@@
$\omega_{c}=\frac{qB_{\text{eff}}}{m}$ (7)
which depends on mass m, charge q, and relevant magnetic field B leading to species separation, whereas the rigid rotor rotation relies on centrifugal potential energy acceleration to accomplish separation. Both modes of separation require a longitudinal magnetic field directed along the z axis of the rotating cylinder.
Equation 7 provides a way of extracting local magnetic fields, based on individual cyclotron rotation rates, as obtained from measured mass separation ratios. $B_{\text{eff}}$ represents an effective magnetic field extracted from the data as further extensively discussed in Section 7 below.
Shown below in Figure 6 are the results of using Eq. 7, for charge state 1 (which is the most abundant ion), to calculate the magnetic fields experienced by the alloy elements and the isotopes using their extracted rotation rates from Table 1 above. The numerical average taken from all the data points in Fig.6 provides an effective field value of 19.3 Megagauss [MG] within the ablation cylinder. This effective B field will be discussed further in Section 7 below where Ion Bernstein Wave contributions are presented.
Figure 6: Gaussian (isotopes) and quadratic (alloy) fit to cyclotron rotation rates and magnetic field as function of plasma radius and observation angle.
@@
Given this information about rotation rates, we can estimate the number of rotations an ion experiences in traveling the 125 $\mu$ length of our idealized centrifuge cylinder. We show, in Table 2 below, the distance traveled per rotation and the number of rotations achieved by an 0.2 keV Ni isotope ion traveling the length of the cylinder. It is seen in Figure 4 that a charge state one ion has this characteristic energy and represents the bulk of the ionic component in the plasma. The results in Table 2 are computed using separation ratios from Figure 3 and rotation rates from Eq.3. In Fig.3 the total enrichment ratios (for all charge states at a given angle) is shown as circles connected by a solid line. The distance traveled per rotation scales as $E^{1/2}$ for the higher energy ions with a commensurate decrease in number of predicted rotations. At 60 degrees (which has the slowest rotation rate), one full rotation would be accomplished by a 3.2 keV ion traveling the 125 microns. This represents a charge state between 4 and 5 according to Figure 9.